# An asymptotic stability proof and a port-Hamiltonian physics-informed neural network approach to chaotic synchronization

### v2

@@ Abstract
We study chaotic synchronization in a five-dimensional Hindmarsh–Rose neuron model augmented with electromagnetic induction and a switchable memristive autapse. For two diffusively coupled neurons, we derive the linearized error dynamics and prove global asymptotic stability of the synchronization manifold using a quadratic Lyapunov function. Verifiable sufficient conditions follow from Sylvester’s criterion, and convergence is established using Barbalat’s lemma. A Helmholtz decomposition splits the error field into conservative and dissipative parts, yielding a closed-form synchronization energy and its dissipation law, which quantify the energetic cost of synchrony. Numerical simulations confirm complete synchronization, monotone decay of the synchronization energy, and close agreement between Lyapunov and Hamiltonian diagnostics across parameter sweeps. Building on these results, we introduce a port-Hamiltonian physics-informed neural network that embeds the conservative/dissipative structure through tailored losses and structural priors. The learned Hamiltonian and energy-rate match analytical benchmarks, providing an energy-aware, data-driven template for modeling and control of nonlinear neuronal synchronization.
@@
biological neurons, synchronization, Lyapunov function, synchronization energy, Hamilton function, physics-informed neural networks, port-Hamiltonian neural networks
@@ 1 Introduction
Among the various notions of synchrony—such as phase synchronization [7, 8, 9, 10, 11], cluster synchronization [12], phase-synchronized clusters [13, 14], generalized synchronization [15, 16], reduced-order synchronization [17, 18, 19, 20], and increased-order synchronization [21, 22]—we focus here on complete synchronization (CS) of same-order chaotic systems [23, 24]. In this regime, the trajectories of two coupled systems converge so that $\lvert x_{1}(t)-x_{2}(t)\rvert\to 0$ as $t\to\infty$ [6, 23, 25, 26]. CS provides the most direct measure of coherence and a tractable setting for Lyapunov stability analysis. Although other types—such as phase synchronization, characterized by $2\pi$-locked phases with uncorrelated amplitudes—play essential roles in distributed brain dynamics, we emphasize CS as a foundation for understanding the energetic and structural basis of synchronization in nonlinear neurons.
@@
Synchronization has been widely analyzed in neural models of varying complexity [27, 28, 29]. Interactions between excitatory and inhibitory populations regulate both phase and anti-phase locking [30], while coupling strength and connection probability influence the onset of spike or bursting synchronization [31]. Historically, these phenomena have been studied using classical dynamical-systems theory, providing analytical and numerical tools to characterize coherent activity. Within this framework, the present work investigates CS between two electromagnetically and memristively coupled Hindmarsh–Rose (HR) neurons. We provide a Lyapunov-based proof of asymptotic stability, corroborated by numerical simulations, and introduce a physics-informed machine-learning framework that learns the associated energy landscape directly from data.
@@
In recent years, the physics-informed neural network (PINN) framework has become a powerful approach for solving forward and inverse problems governed by differential equations [32, 33, 34]. The key idea is to embed governing physical laws—typically expressed as ordinary or partial differential equations—directly into the learning objective. By penalizing equation residuals during training, the model enforces physical consistency and improves generalization beyond observed data [34]. For example, [33] demonstrated that causality-aware loss functions enable stable learning of chaotic and turbulent systems such as the Lorenz attractor and the Navier–Stokes equations.
@@
A complementary development is the Hamiltonian neural network (HNN) [35], which learns the underlying energy function of a dynamical system. An HNN represents the dynamics using a learned Hamiltonian $H_{\theta}(x)$ together with a fixed skew-symmetric interconnection matrix enforcing energy conservation. This enables unsupervised identification of conservative dynamics from trajectory data. Extensions such as [36] and [37] incorporated structural priors, coordinate transformations, and learned symplectic forms to improve identifiability and data efficiency. Nonetheless, classical HNNs apply only to conservative systems and cannot represent dissipation or environmental interactions, which are fundamental in neuronal dynamics.
@@
Beyond general dynamical-systems modeling, these neural frameworks have growing relevance in computational neuroscience, where nonlinear oscillators such as FitzHugh–Nagumo or HR neurons display multistability, bursting, and chaotic synchronization. Physics-informed and Hamiltonian-based learning methods offer an interpretable bridge between data-driven modeling and biophysical realism. By incorporating principles such as energy conservation and dissipation, they help illuminate the energetic mechanisms underlying neuronal synchrony—a key determinant of coherent neural activity and information processing. Unifying Hamiltonian structure with physics-informed regularization thus provides a principled and biologically meaningful way to study the energetics of synchronization and decoherence in complex neural systems.
Motivated by these challenges, we retain the structural advantages of the port–Hamiltonian formulation while regularizing learning through PINN-style residual constraints. To our knowledge, no prior work has unified port–Hamiltonian factorization with physics-informed residual training in a single framework. We therefore propose the port–Hamiltonian physics-informed neural network (pH–PINN), which combines the interpretability of energy-based modeling with the consistency of residual-driven training. Specifically, the pH–PINN enforces the governing differential equations and uses a Helmholtz decomposition of the vector field to associate the divergence-free (energy-conserving) component with the Hamiltonian interconnection flow and the curl-free (dissipative) component with the non-conservative flow. This dual constraint improves identifiability, prevents degenerate energy landscapes, and alleviates the need for long rollout windows—making it well-suited for studying the energetic and stability aspects of synchronization in chaotic systems. As shown here, the pH–PINN accurately recovers both the Hamiltonian function and its time derivative, bridging analytical dynamical systems theory with data-driven discovery.
The paper is structured as follows: Section 2 introduces the 5D HR neuron model and its dynamics, illustrated through bifurcation diagrams. Section 3 establishes asymptotic stability of the synchronization manifold via Lyapunov analysis. Section 4 presents numerical simulations that corroborate the theoretical results. Section 5 develops the pH–PINN framework for data-driven learning of synchronization energetics. Finally, Section 6 summarizes the findings and discusses potential extensions.
@@ 2 Mathematical model and dynamics
We investigate a paradigmatic neuron model of established biological relevance—the HR model [44, 45]—extended to incorporate additional biophysical mechanisms. Specifically, we consider the effects of electromagnetic induction [46, 47] and an autaptic feedback current governed by a switchable memristive synapse, resulting into a new five-dimensional formulation of the HR neuron model is given by [48]:
$\left\{\begin{aligned} \dot{x}&=y-ax^{3}+bx^{2}+k(h+fu^{2})x+\rho\phi x+I,\\ \dot{y}&=c-dx^{2}-y,\\ \dot{z}&=r\left[s(x+x_{0})-z\right],\\ \dot{u}&=-u+m\left(|u+1|-|u-1|\right)+x,\\ \dot{\phi}&=x-q\phi.\end{aligned}\right$. (1)
where $x$ and $y$ represent the membrane potential and the fast recovery variable, respectively, while $z$ accounts for the slow adaptation current. The variable $u$ denotes the internal state of the switchable memristor governing the autaptic feedback current, and $\phi$ represents the magnetic flux associated with electromagnetic induction. The parameters $a,b,c,d,r,s,k,h,f,m,q,\rho,x_{0},$ and $I$ are real constants that control the intrinsic dynamics, coupling, and feedback strength of the system.
@@
Due to the strong nonlinearities in Eq. (1), the system was analyzed via high-precision numerical integration. Simulations were carried out in Python using JAX [50] and Diffrax [51] with the adaptive Tsitouras 5/4 Runge–Kutta method (Tsit5) [52]. Computations employed 64-bit precision with tolerances of $10^{-8}$ (relative) and $10^{-10}$ (absolute). Each trajectory was integrated for 1000 time units from $\mathbf{x}_{0}=(0,0,0,1,0)^{\top}$, discarding the first 500 units to remove transients. Bifurcation diagrams and Lyapunov spectra were obtained from the post-transient local maxima of $x(t)$, and Lyapunov exponents were computed using the Benettin algorithm with periodic orthonormalization. The resulting bifurcation structures are shown in Figs. 24.
@@
Figure 2 shows the bifurcation diagrams and the two largest Lyapunov exponents versus $k$ for fixed $\rho=0.7$ and different $m$. As $k$ increases, the system exhibits a period-doubling route to chaos, while smaller $m$ values suppress chaotic oscillations.
@@
Figure 1: Chaotic time series of the $x$ variable (a) and corresponding phase portrait in the xyx-y plane (b) with $k=0.87$, $m=0.5$ and $\rho=0.7$.
Figure 2: Bifurcation diagrams and the two largest Lyapunov exponents versus $k$ at $\rho=0.7$ for different values of $m$: (a) $m=1$, (b) $m=0.5$, (c) $m=0.25$.
Figure 3: Bifurcation diagrams and the two largest Lyapunov exponents versus $\rho$ at $k=0.87$ for different values of $m$: (a) $m=1$, (b) $m=0.5$, (c) $m=0.25$.
@@
3 Stable chaotic synchronization energy: A dynamical systems approach
In this section, we investigate the stability of CS between two diffusively coupled chaotic 5D HR neurons governed by Eq. (1), together with the associated synchronization energy. The analysis is conducted within the framework of dynamical systems theory, employing Lyapunov stability analysis [53] to evaluate the onset and robustness of synchronization, and Helmholtz’s theorem [54] to characterize the energetic aspects of the synchronized state. Particular attention is devoted to the asymptotic stability of the synchronization manifold and to the behavior of the synchronization energy under variations of the key system parameters.
3.1 Asymptotic stability proof of the synchronization manifold
@@
$\left\{\begin{aligned} \dot{x}_{i}&=y_{i}-ax_{i}^{3}+bx_{i}^{2}+k(h+fu_{i}^{2})x_{i}+\rho\phi_{i}x_{i}+I_{i}+g_{e}\sum_{j=1}^{2}\xi_{ij}(x_{j}-x_{i}),\\ \dot{y}_{i}&=c-dx_{i}^{2}-y_{i},\\ \dot{z}_{i}&=r\left[s(x_{i}+x_{0})-z_{i}\right],\\ \dot{u}_{i}&=-u_{i}+m\left(|u_{i}+1|-|u_{i}-1|\right)+x_{i},\\ \dot{\phi}_{i}&=x_{i}-q\phi_{i},\end{aligned}\right$. (2)
where $g_{e}$ denotes the electrical (gap-junction) coupling strength that determines the intensity of interaction between neurons. The coupling topology is defined by $\xi_{ij}=1$ for $i\neq j$, indicating mutual coupling between neurons $i$ and $j$, and $\xi_{ij}=0$ otherwise. For the two-neuron system considered here, $\xi_{12}=\xi_{21}=1$. This formulation provides the mathematical basis for studying how the interplay between intrinsic neuronal dynamics and diffusive coupling gives rise to collective synchronization.
Definition 3.1.
Two chaotic systems $x(t),y(t)\in\mathbb{R}^{n}$ are said to achieve CS if there exist initial conditions $x(t_{0})=x_{0}$ and $y(t_{0})=y_{0}$ such that
$\lim_{t\to\infty}\|x(t)-y(t)\|=0.$ (3)
@@
where $\mathbf{x}_{s}=(x_{s},y_{s},z_{s},u_{s},\phi_{s})^{\mathsf{T}}$ evolves according to the uncoupled dynamics given in Eq. (1). To investigate the global asymptotic stability of $\mathcal{M}$, we introduce the synchronization error system via
@@
By taking the difference between the equations of the two coupled neurons, one arrives at the corresponding error dynamics. Discarding higher-order nonlinear contributions in the error variables ($e_{x}^{2}$, $e_{x}^{3}$, $e_{u}^{2}$, $e_{x}e_{\phi}$), we obtain the linearized error system that describes the most significant transverse deviations from $\mathcal{M}$:
@@
Theorem 3.2.
The coupled chaotic systems in Eq. (2) achieve CS if the five ordered principal minors in Eq. (15) are all sufficiently positive.
@@
$V(e_{x},e_{y},e_{z},e_{u},e_{\phi})=\tfrac{1}{2}\left(e_{x}^{2}+e_{y}^{2}+e_{z}^{2}+e_{u}^{2}+e_{\phi}^{2}\right).$ (7)
@@
Since the neurons evolve on bounded chaotic attractors, we can assert the existence of a constant $J>0$ such that

### v3

# A Lyapunov stability proof and a port-Hamiltonian physics-informed neural network for chaotic synchronization in memristive neurons

@@ Abstract
We study chaotic synchronization in a 5D Hindmarsh–Rose neuron model augmented with electromagnetic induction and a switchable memristive autapse. For two diffusively coupled identical neurons, we derive the transverse error dynamical system and analyze local synchronization via the linearized error system around the synchronization manifold. A quadratic Lyapunov function yields explicit sufficient conditions for (i) asymptotic stability when the memristive switching remains dissipative and (ii) practical stability with an explicit ultimate bound under non-dissipative switching. We complement this with a Hamiltonian-based viewpoint: a Helmholtz decomposition of the linearized error vector field provides a closed-form synchronization Hamiltonian and its rate identity. Numerical simulations corroborate convergence or ultimate boundedness of the synchronization errors and an overall decay of the synchronization Hamiltonian and its instantaneous rate toward zero after transients, and show consistent trends between Lyapunov- and Hamiltonian-based diagnostics across parameters. Finally, we propose the first port-Hamiltonian physics-informed neural network (pH-PINN) that learns this synchronization Hamiltonian and its rate from data while preserving conservative/dissipative structure, achieving close agreement with the analytical expressions.
@@
memristive neurons , chaos , synchronization, Lyapunov function, Hamiltonian function , port-Hamiltonian neural networks, physics-informed neural networks
@@ 1 Introduction
Among the various notions of synchrony—such as phase synchronization [7, 8, 9, 10, 11], cluster synchronization [12], phase-synchronized clusters [13, 14], generalized synchronization [15, 16], reduced-order synchronization [17, 18, 19, 20], and increased-order synchronization [21, 22]—we focus here on complete synchronization of same-order chaotic systems [23, 24]. In this regime, the trajectories of two coupled systems converge so that $\lvert x_{1}(t)-x_{2}(t)\rvert\to 0$ as $t\to\infty$ [6, 23, 25, 26]. Complete synchronization provides the most direct measure of coherence and a tractable setting for Lyapunov stability analysis. Although other types—such as phase synchronization, characterized by $2\pi$-locked phases with uncorrelated amplitudes—play essential roles in distributed brain dynamics, we emphasize complete synchronization as a foundation for understanding the energetic and structural basis of synchronization in nonlinear neurons.
@@
Synchronization has been widely analyzed in neural models of varying complexity [27, 28, 29]. Interactions between excitatory and inhibitory populations regulate both phase and anti-phase locking [30], while coupling strength and connection probability influence the onset of spike or bursting synchronization [31]. Historically, these phenomena have been studied using classical dynamical systems theory, providing analytical and numerical tools to characterize coherent activity. Within this framework, the present work investigates complete synchronization between two electromagnetically and memristively coupled Hindmarsh–Rose (HR) neurons. We provide a Lyapunov-based proof of asymptotic and practical stability, corroborated by numerical simulations, and introduce a physics-informed machine-learning framework that learns the associated Hamiltonian landscape directly from data, while respecting the conservative/dissipative structure of the synchronization Hamiltonian.
@@
In recent years, the physics-informed neural network (PINN) framework has become a powerful approach for solving forward and inverse problems governed by differential equations [32, 33, 34]. The key idea is to embed governing physical laws—typically expressed as ordinary or partial differential equations—directly into the learning objective. By penalizing equation residuals during training, the model enforces physical consistency and improves generalization beyond observed data [34]. For example, [33] demonstrated that causality-aware loss functions enable stable learning of chaotic and turbulent systems such as the Lorenz attractor and the Navier–Stokes equations. Along related lines, Savaliya and Yamakou [35] used a noise-augmented PINN surrogate to model and predict self-induced stochastic resonance in the stochastic FitzHugh–Nagumo neuron by embedding the governing SDE residuals together with Kramers escape-time (barrier-matching) constraints directly into the training loss.
@@
A complementary development is the Hamiltonian neural network (HNN) [36], which learns the underlying Hamiltonian function of a dynamical system. An HNN represents the dynamics using a learned Hamiltonian $H_{\theta}(x)$ together with a fixed skew-symmetric interconnection matrix enforcing Hamiltonian conservation. This enables unsupervised identification of conservative dynamics from trajectory data. Extensions such as [37] and [38] incorporated structural priors, coordinate transformations, and learned symplectic forms to improve identifiability and data efficiency. Nonetheless, classical HNNs apply only to conservative systems and cannot represent dissipation or environmental interactions, which are fundamental in neuronal dynamics.
@@
Beyond general dynamical systems modeling, these neural frameworks have growing relevance in computational neuroscience, where nonlinear oscillators such as FitzHugh–Nagumo or HR neurons display multistability, bursting, and chaotic synchronization. Physics-informed and Hamiltonian-based learning methods offer an interpretable bridge between data-driven modeling and biophysical realism. By incorporating principles such as energy conservation and dissipation, they help illuminate the energetic mechanisms underlying neuronal synchrony—a key determinant of coherent neural activity and information processing. Unifying Hamiltonian structure with physics-informed regularization thus provides a principled and biologically meaningful way to study the energetics of synchronization and decoherence in complex neural systems.
Motivated by these challenges, we retain the structural advantages of the port–Hamiltonian formulation while regularizing learning through PINN-style residual constraints. To our knowledge, no prior work has unified port–Hamiltonian factorization with physics-informed residual training in a single framework. We therefore propose the port–Hamiltonian physics-informed neural network (pH–PINN), which combines the interpretability of Hamiltonian -based modeling with the consistency of residual-driven training. Specifically, the pH–PINN enforces the governing differential equations and uses a Helmholtz decomposition of the vector field to associate the divergence-free ( Hamiltonian -conserving) component with the Hamiltonian interconnection flow and the curl-free (dissipative) component with the non-conservative flow. This dual constraint improves identifiability, prevents degenerate Hamiltonian landscapes, and alleviates the need for long rollout windows—making it well-suited for studying the energetic and stability aspects of synchronization in chaotic systems. As shown here, the pH–PINN accurately recovers both the Hamiltonian function and its time derivative, bridging analytical dynamical systems theory with data-driven discovery.
The paper is structured as follows: Section 2 introduces the 5D HR neuron model and its dynamics, illustrated through bifurcation diagrams. Section 3 establishes asymptotic and practical stability of the synchronization manifold via Lyapunov analysis. Section 4 presents numerical simulations that corroborate the theoretical results. Section 5 develops the pH–PINN framework for data-driven learning of the synchronization Hamiltonian and its rate identity . Finally, Section 6 summarizes the findings and discusses potential extensions.
@@ 2 Mathematical model and dynamics
We investigate a paradigmatic neuron model of established biological relevance—the HR model [45, 46]—extended to incorporate additional biophysical mechanisms. Specifically, we consider the effects of electromagnetic induction [47, 48] and an autaptic feedback current governed by a switchable memristive synapse, resulting in a new five-dimensional formulation of the HR neuron model [49] is given by
$\left\{\begin{aligned} \dot{x}&=y-ax^{3}+bx^{2}+k(h+fu^{2})x+\rho\phi x+I,\\ \dot{y}&=c-dx^{2}-y,\\ \dot{z}&=r\left[s(x+{\color[rgb]{0,0,0}x_{r}})-z\right],\\ \dot{u}&=-u+m\left(|u+1|-|u-1|\right)+x,\\ \dot{\phi}&=x-q\phi.\end{aligned}\right$. (1)
where $x$ and $y$ represent the membrane potential and the fast recovery variable, respectively, while $z$ accounts for the slow adaptation current. The variable $u$ denotes the internal state of the switchable memristor governing the autaptic feedback current, and $\phi$ represents the magnetic flux associated with electromagnetic induction. The parameters $a,b,c,d,r,s,k,h,f,m,q,\rho,{\color[rgb]{0,0,0}x_{r}},$ and $I$ are real constants that control the intrinsic dynamics, coupling, and feedback strength of the system.
@@
These parameters play distinct roles associated with the two added biophysical mechanisms. Electromagnetic induction is captured by the magnetic flux variable $\phi$, whose relaxation rate is controlled by $q$ in $\dot{\phi}=x-q\phi$, and whose feedback onto the membrane potential is governed by the coupling coefficient $\rho$ through the term $\rho\,\phi\,x$ [50]. Memristive/autaptic feedback is captured by the internal state $u$, whose switching nonlinearity is controlled by $m\in[0,1]$ via m(|u+1||u1|)m(|u+1|-|u-1|), and whose modulation of the membrane potential is weighted by the gain $k$ and parameters $h,f$ through the term $k(h+fu^{2})x$ [49]. Accordingly, varying $\rho$ and $q$ primarily tunes the induction pathway, whereas varying $m$, $k$, $h$, and $f$ tunes the strength and nonlinearity of the memristive/autaptic pathway. Variations in these parameters lead to transitions among equilibrium states, periodic oscillations, and chaotic attractors, revealing a rich bifurcation structure underlying the model’s nonlinear behavior.
Due to the strong nonlinearities in Eq. (1), the system was analyzed via high-precision numerical integration. Simulations were carried out in Python using JAX [51] and Diffrax [52] with the adaptive Tsitouras 5/4 Runge–Kutta method (Tsit5) [53]. Computations employed 64-bit precision with tolerances of $10^{-8}$ (relative) and $10^{-10}$ (absolute). Each trajectory was integrated for 1000 time units from the initial state $\mathbf{X}(0)=(0,0,0,1,0)^{\mathsf{T}}$ , discarding the first 500 units to remove transients. Bifurcation diagrams and Lyapunov spectra were obtained from the post-transient local maxima of $x(t)$, and Lyapunov exponents were computed using the Benettin algorithm with periodic orthonormalization. The resulting bifurcation structures are shown in Figs. 24.
@@
Figure 2 shows the bifurcation diagrams and the two largest Lyapunov exponents (LEs) versus $k$ for fixed $\rho=0.7$ and different $m$. As $k$ increases, the system exhibits a period-doubling route to chaos, while smaller $m$ values suppress chaotic oscillations.
@@
Figure 1: Chaotic time series of the $x$ variable (a) and corresponding phase portrait in the xyx-y plane (b) with $k=0.87$, $m=0.5$, and $\rho=0.7$.
Figure 2: Bifurcation diagrams and the two largest Lyapunov exponents (LEs) versus $k$ at $\rho=0.7$ for different values of $m$: (a) $m=1$, (b) $m=0.5$, (c) $m=0.25$.
Figure 3: Bifurcation diagrams and the two largest Lyapunov exponents (LEs) versus $\rho$ at $k=0.87$ for different values of $m$: (a) $m=1$, (b) $m=0.5$, (c) $m=0.25$.
@@
3 Synchronization and its Hamiltonian: A dynamical systems approach
In this section, we investigate the stability of complete synchronization between two diffusively coupled chaotic 5D HR neurons governed by Eq. (1), together with the associated synchronization Hamiltonian . The analysis is conducted within the framework of dynamical systems theory, employing Lyapunov stability analysis [54] to evaluate the onset and robustness of synchronization, and Helmholtz’s theorem [55] to characterize the energetic aspects of the synchronized state. Particular attention is devoted to the asymptotic and practical stability of the synchronization manifold associated with the linearized error system and to the behavior of its synchronization Hamiltonian under variations of the key system parameters.
3.1 Asymptotic and practical stability in the linearized error dynamical system
@@
$\left\{\begin{aligned} \dot{x}_{i}&=y_{i}-ax_{i}^{3}+bx_{i}^{2}+k(h+fu_{i}^{2})x_{i}+\rho\phi_{i}x_{i}+I_{i}+g_{e}\sum_{j=1}^{2}\xi_{ij}(x_{j}-x_{i}),\\ \dot{y}_{i}&=c-dx_{i}^{2}-y_{i},\\ \dot{z}_{i}&=r\left[s(x_{i}+{\color[rgb]{0,0,0}x_{r}})-z_{i}\right],\\ \dot{u}_{i}&=-u_{i}+m\left(|u_{i}+1|-|u_{i}-1|\right)+x_{i},\\ \dot{\phi}_{i}&=x_{i}-q\phi_{i},\end{aligned}\right$. (2)
where $g_{e}$ denotes the electrical (gap-junction) coupling strength that determines the intensity of interaction between neurons; throughout this paper we fix $g_{e}=0.62$ unless stated otherwise. The coupling topology is defined by $\xi_{ij}=1$ for $i\neq j$, indicating mutual coupling between neurons $i$ and $j$, and $\xi_{ij}=0$ otherwise. For the two-neuron system considered here, $\xi_{12}=\xi_{21}=1$. This formulation provides the mathematical basis for studying how the interplay between intrinsic neuronal dynamics and diffusive coupling gives rise to collective synchronization.
Definition 3.1 (Complete synchronization).
Two dynamical systems with states $x(t),y(t)\in\mathbb{R}^{n}$ are said to achieve complete synchronization if there exist initial states $x(t_{0})=x_{0}$ and $y(t_{0})=y_{0}$ such that
$\lim_{t\to\infty}\|x(t)-y(t)\|=0.$
Definition 3.2 (Practical ($\delta$-)synchronization).
Two dynamical systems with states $x(t),y(t)\in\mathbb{R}^{n}$ are said to achieve practical (or $\delta$-) synchronization if there exist initial states $x(t_{0})=x_{0}$, $y(t_{0})=y_{0}$ and a tolerance $\delta>0$ such that
$\limsup_{t\to\infty}\|x(t)-y(t)\|\leq\delta.$
@@
where $\mathbf{X}_{s}=(x_{s},y_{s},z_{s},u_{s},\phi_{s})^{\mathsf{T}}$ evolves according to the uncoupled dynamics given in Eq. (1). To investigate the stability of $\mathcal{M}$, we introduce the synchronization error system via
@@
By taking the difference between the equations of the two coupled neurons, one arrives at the corresponding full nonlinear error dynamical system given by
@@
Discarding second and higher-order nonlinear contributions in the errors ($e_{x}^{2}$, $e_{u}^{2}$, $e_{x}e_{\phi}$, $e_{x}e_{u}$, $e_{x}^{3}$, $e_{x}e_{u}^{2}$), we obtain the linearized error dynamical system that describes the most significant transverse deviations from $\mathcal{M}$ as follow:
[…]
Remark 1.
Equation (6) represents the linearized transverse error dynamics around the synchronization manifold $\mathcal{M}$. The Lyapunov/Barbalat analysis below therefore establishes asymptotic and practical convergence for this linearized error system (local transverse stability in the usual sense). A fully global nonlinear synchronization proof hinges explicit bounds on the neglected nonlinear remainder terms, and obtaining such semi-global or global estimates is beyond the scope of the current work. Nevertheless, alongside the numerics of the linearized error system (Eq. (6)) synchronization in Section 4, we also present the synchronization of the full nonlinear error system in Eq. (5) as numerical evidence (Fig. 5(a1)-(e1)).
Theorem 3.3 (Asymptotic stability under dissipative memristive contribution).
Consider the linearized transverse error dynamics Eq. (6) along a bounded reference trajectory $(x_{1}(t),u_{1}(t),\phi_{1}(t))$ of Eq. (2), and assume there exists $J>0$ such that $|x_{1}(t)|,|u_{1}(t)|,|\phi_{1}(t)|\leq J$ for all $t\geq 0$. Let $\mathbf{M}$ be the diagonal (hence symmetric) matrix defined in Eq. (18), and assume that all its diagonal entries are strictly positive so that $\mathbf{M}\succ 0$.
Assume in addition that along the reference trajectory, the memristive contribution is dissipative in the sense that
$\widehat{\Psi}(e_{u};u_{1}(t),u_{2}(t),m,J)\leq-e_{u}^{2},\qquad\forall t\geq 0,$
for all $e_{u}$, i.e., the switching remains in the dissipative Cases $2,3,5,8$ of Eq. (14).
Then the fixed point $\mathbf{e}\equiv 0$ of Eq. (6) is asymptotically stable, and in particular
$\lim_{t\to\infty}\|\mathbf{e}(t)\|=0.$
@@
$V(e_{x},e_{y},e_{z},e_{u},e_{\phi})=\tfrac{1}{2}\left(e_{x}^{2}+e_{y}^{2}+e_{z}^{2}+e_{u}^{2}+e_{\phi}^{2}\right)\geq 0.$ (7)
@@
We assume solutions of the HR neuron system remain bounded (as observed numerically and consistent with the existence of a chaotic attractor). Hence there exists a constant $J>0$ such that
@@
$\Psi(e_{u};u_{1},u_{2},m,J)=\begin{cases}|2m-1|e_{u}^{2}+2|m||J-1||e_{u}|,&\text{if }u_{1}\geq 1,\;-1<u_{2}<1,\\ -e_{u}^{2}-4|m||e_{u}|,&\text{if }u_{1}\geq 1,\;u_{2}\leq-1,\\ -e_{u}^{2}-2|m||J-1||e_{u}|,&\text{if }-1<u_{1}<1,\;u_{2}\geq 1,\\ |2m-1|e_{u}^{2},&\text{if }-1<u_{1}<1,\;-1<u_{2}<1,\\ -e_{u}^{2}-2|m|(J+1)|e_{u}|,&\text{if }-1<u_{1}<1,\;u_{2}\leq-1,\\ -e_{u}^{2}+4|m||e_{u}|,&\text{if }u_{1}\leq-1,\;u_{2}\geq 1,\\ |2m-1|e_{u}^{2}+2|m|(J+1)|e_{u}|,&\text{if }u_{1}\leq-1,\;-1<u_{2}<1,\\ -e_{u}^{2},&\text{otherwise.}\end{cases}$ (11)
Next, we make explicit the inequality step used to pass from the componentwise estimate Eq. (10) to a compact quadratic-form bound. We invoke Young’s inequalities in the following forms: (1) for any $\alpha,\beta\geq 0$, $2\alpha\beta\leq\alpha^{2}+\beta^{2}$ and (2) for any $\alpha\in\mathbb{R}$, any $c\geq 0$, and any $\varepsilon>0$, $c|\alpha|\leq\frac{\varepsilon}{2}\alpha^{2}+\frac{c^{2}}{2\varepsilon}.$ Applying (1) to each mixed product term in Eq. (10) yields
$\left\{\begin{aligned} (1+2|d|J)\,|e_{x}|\,|e_{y}|&\leq\frac{1}{2}\,(1+2|d|J)\,\Big(e_{x}^{2}+e_{y}^{2}\Big),\\ |rs|\,|e_{x}|\,|e_{z}|&\leq\frac{1}{2}\,|rs|\,\Big(e_{x}^{2}+e_{z}^{2}\Big),\\ (2|kf|J^{2}+1)\,|e_{x}|\,|e_{u}|&\leq\frac{1}{2}\,(2|kf|J^{2}+1)\,\Big(e_{x}^{2}+e_{u}^{2}\Big),\\ (|\rho|J+1)\,|e_{x}|\,|e_{\phi}|&\leq\frac{1}{2}\,(|\rho|J+1)\,\Big(e_{x}^{2}+e_{\phi}^{2}\Big),\end{aligned}\right$. (12)
where $\mathbf{e}=\left(e_{x},e_{y},e_{z},e_{u},e_{\phi}\right)^{\top}$ denotes the signed error vector components. Applying (2) to the linear $|e_{u}|$-terms arising from the memristive contribution in Eq. (11) (and, in Cases $2,3,5$, using the sharper trivial bound $\Psi\leq-e_{u}^{2}$ since the corresponding linear term is nonpositive) yields
$\Psi(e_{u};u_{1},u_{2},m,J)\leq\widehat{\Psi}(e_{u};u_{1},u_{2},m,J),$ (13)
where $\widehat{\Psi}$ is given in Eq. (14).